OpenStudy (anonymous):

Medal and Fan Evaluate the expression: [(8 × 4) − 10]2 + 9

OpenStudy (anonymous):

First work out everything in the brackets [ ]

OpenStudy (anonymous):

(8 x 4) - 10

OpenStudy (anonymous):

22

OpenStudy (anonymous):

(22) 2 + 9

OpenStudy (anonymous):

53

OpenStudy (anonymous):

Yeah exactly

OpenStudy (anonymous):

no its 493

OpenStudy (anonymous):

because its 22 to the 2nd power

OpenStudy (anonymous):

no, if it was 22 to the second power it would look like \[(22)^2\]

OpenStudy (anonymous):

this is (22)2, which means multiply 22 times 2

OpenStudy (anonymous):

it is i just copy and pasted it

OpenStudy (anonymous):

1. Work With Brackets "8*4 2. Then Subtract By 10 3. Answer * 2 4 Add 9

OpenStudy (anonymous):

its 22^2 i just copy and pasted it and it went wrong

OpenStudy (anonymous):

Oh okay, then yeah, it is 493

OpenStudy (kendricklamar2014):

Let go Step by Step: Step 1. 8 * 3 = 32 (32 - 10)^2 + 9 Step 2. 32 - 10^2 = 22^2 = 22^2 + 9 Step 3. 22^2 = 484 = 484 + 9 = 493 Answer: 493

OpenStudy (kendricklamar2014):

Hope it helps :)

OpenStudy (anonymous):

yea i got it already thanks

OpenStudy (kendricklamar2014):

No Problem